Session: Feb 1999 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Sat Feb 27 01:04:43 MET 1999) Success
No known problem

Torun (K. Borkowski at Sun Feb 28 15:32:23 MET 1999) Success
No known problems.

Shanghai (Huang Xinyong at Mon Mar 1 11:57:04 MET 1999) Success
No problems during the observation. It is raining.

Westerbork (Tony Foley at Mon Mar 1 12:28:15 MET 1999) Minor failures
Vacuum loss caused loss of data from 19:52 to 20:27UTl; tape gave lots of dirt in vacuum chamber giving problem with sensors

Noto (C.Trigilio at Mon Mar 1 22:52:55 MET 1999) Success

Jodrell (Alastair Gunn at Tue Mar 2 05:41:51 MET 1999) Success
No known problems.

Effelsberg (dgraham at Tue Mar 2 13:40:10 MET 1999) Minor failures
1940-1951 offsrc,2204-2209 frequencies wrong, otherwise OK. Cannot see any lines on autocorr spectrum with realtime system

Onsala (A. Polatidis at Wed Mar 3 00:55:49 MET 1999) Success

Calibration (Jean\-Francois Desmurs at Wed Mar 3 18:32:32 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: gy001c.antab. Read the comments in the file to get further information. (Sh no Tsys values on RCP).